OpenStudy (anonymous):

The displacement (in centimeters) of a particle moving back and forth along a straight line is given by the equation of motion s = 2 sin πt + 3 cos πt, where t is measured in seconds. (Round your answers to two decimal places.) (a) Find the average velocity during each time period. (i) [1, 2] cm/s (ii) [1, 1.1] cm/s (iii) [1, 1.01] cm/s (iv) [1, 1.001] cm/s (b) Estimate the instantaneous velocity of the particle when t = 1. cm/s

OpenStudy (anonymous):

average velocity = ds/dt =d/dt ( 2 sin πt + 3 cos πt) = 2π cos πt - 3πsin πt if u substitute the values of time interval as upper limit and lower limit u will get the average velocity